# TUS-REC2024: A Challenge to Reconstruct 3D Freehand Ultrasound Without External Tracker

### v1

@@ Abstract
Trackerless freehand ultrasound reconstruction aims to reconstruct 3D volumes from sequences of 2D ultrasound images without relying on external tracking systems, offering a low-cost, portable, and widely deployable alternative for volumetric imaging. However, it presents significant challenges, including accurate inter-frame motion estimation, minimisation of drift accumulation over long sequences, and generalisability across scanning protocols. The TUS-REC2024 Challenge was established to benchmark and accelerate progress in trackerless 3D ultrasound reconstruction by providing a publicly available dataset for the first time, along with a baseline model and evaluation framework. The Challenge attracted over 43 registered teams, of which 6 teams submitted 21 valid dockerized solutions. Submitted methods spanned a wide range of algorithmic approaches, including recurrent models, registration-driven volume refinement, attention, and physics-informed models. This paper presents an overview of the Challenge design, summarises the key characteristics of the dataset, provides a concise literature review, introduces the technical details of the underlying methodology working with tracked freehand ultrasound data, and offers a comparative analysis of submitted methods across multiple evaluation metrics. The results highlight both the progress and current limitations of state-of-the-art approaches in this domain, and inform directions for future research. The data, evaluation code, and baseline are publicly available to facilitate ongoing development and reproducibility. As a live and evolving benchmark, this Challenge is designed to be continuously developed and improved. The Challenge was held at MICCAI 2024 and will be organised again at MICCAI 2025, reflecting its growing impact and the sustained commitment to advancing this field.
@@
Trackerless , Freehand , Ultrasound , 3D reconstruction , Spatial transformation estimation , MICCAI 2024 challenge , TUS-REC2024
@@ 1 Introduction
Ultrasound imaging remains a cost-effective, non-invasive modality with real-time capabilities, making it a valuable tool across a wide range of clinical applications. Although the underlying data is inherently three-dimensional (3D), standard ultrasound typically captures single two-dimensional (2D) frame without spatial localisation across frames. This poses challenges for applications requiring accurate volumetric information, such as biometric quantification, image registration, and 3D visualisation. While expert clinicians can often infer 3D structure mentally or through standardised acquisition protocols (e.g., standard planes), the absence of inter-frame positional data limits reproducibility and the integration of ultrasound into advanced image analysis workflows.
@@
Ongoing work seeks to address this limitation by using 3D ultrasound probes to enable 3D reconstruction. 3D ultrasound probes is capable of acquiring volumetric data directly, using dedicated mechanical probe or 2D array transducers. While these probes provide valuable 3D imaging capabilities and offer flexible scanning trajectories, their higher cost and limited availability may restrict their use in some clinical settings, such as low-resource environments, point-of-care scenarios, or mobile and emergency units where portability and affordability are critical.
@@
A key advantage of freehand 2D ultrasound is its widespread availability and long-standing integration into clinical workflows. It has been used for decades across a broad range of applications, and clinicians are highly familiar with both its handling and interpretation. Building on this established foundation, tracker-based freehand ultrasound reconstruction techniques have been introduced to enable the generation of 3D anatomical representations. These methods aim to enhance conventional 2D ultrasound by incorporating spatial information from external tracking systems, such as optical or electromagnetic (EM) trackers. This enables conventional 2D ultrasound probes to be used for 3D imaging, providing a more flexible and accessible solution in clinical and research applications where dedicated and bulky 3D ultrasound systems may be impractical. However, optical and EM tracking systems present additional challenges in clinical environments. Optical tracking requires a continuous, unobstructed line of sight between the tracker and the camera, although some solutions, such as using multiple cameras, have been proposed to mitigate this limitation. EM tracking remains sensitive to nearby metal objects and electromagnetic interference, which can affect accuracy.
@@
Trackerless freehand ultrasound reconstruction refers to a class of techniques aimed at generating 3D volumetric representations from sequential 2D ultrasound frames, without the use of external tracking systems. Instead, these methods compute the relative spatial transformations among frames using internal data sources, such as image contents and signals from internal sensors. Common approaches include image-based registration, speckle decorrelation, inertial measurement unit (IMU) integration, or learning-based motion estimation such as convolutional neural networks or recurrent models. Additionally, trackerless freehand ultrasound reconstruction may further enhance existing 3D ultrasound systems, rather than serving solely as alternatives.
@@
However, trackless reconstruction remains challenging due to: 1) the difficulty of maintaining accuracy over long sequences of ultrasound frames, where small frame-to-frame errors can accumulate significantly; and 2) the high variability across different tasks and datasets, which complicates the validation and fair comparison of methods. While benchmarking is essential to address this variability, progress has been limited by the scarcity of publicly available datasets, which are critical for both performance evaluation and the development of learning-based approaches.
@@
This application involves practical challenges such as handling both 2D and 3D imaging data, incorporating tracking information, and managing multiple spatial coordinate systems, all of which contribute to a significant barrier for newcomers to the field and may impede broader progress and adoption. Furthermore, comparisons of methods in the existing literature are often conducted on relatively small, private datasets, using a variety of evaluation metrics to assess performance. This variability complicates the comparison of strengths and weaknesses across methods and may lead to biased conclusions, due to dataset characteristics, evaluation metric choices, and inherent differences in the methods’ underlying assumptions. For example, learning-based approaches may assume that training and testing data come from similar distributions, while classical methods may rely on consistently available speckle patterns.
@@
To address these gaps, we present a significant study of trackerless freehand ultrasound reconstruction, aiming to provide researchers with clear technical insights and consistent terminology. This effort is formalised in the form of the TUS-REC2024 Challenge, which is designed to foster both algorithmic innovation and practical clinical applicability by promoting reproducibility, benchmarking, and methodological transparency. TUS-REC2024 Challenge provides a large-scale in vivo ultrasound dataset, consisting of scans from both the left and right forearms of 85 volunteers (2,040 scans, 1,025,448 frames in total), acquired using a time-synchronised optical tracking system. We aim to conduct a comprehensive comparison among methods, evaluating their strengths and weaknesses on a common, large-scale dataset, using a consistent set of carefully-defined performance metrics. This approach will ensure a more objective and transparent assessment of the methods’ relative efficacy, and more importantly, to drive the development of new techniques for freehand ultrasound reconstruction.
@@
The significance of this Challenge lies in three key contributions. First, it establishes a rigorous and standardised benchmarking framework for trackerless freehand ultrasound reconstruction, advancing the development of novel algorithms and promoting objective performance evaluation through withheld test data and unified assessment metrics. Second, it provides the necessary infrastructure to support this benchmarking effort, including the public release of a comprehensive freehand ultrasound dataset, the largest publicly available dataset to date in the field, alongside detailed preliminary materials and accompanying code that describe the end-to-end pipeline for trackerless reconstruction. Third, beyond the outcomes of the Challenge itself, this summary paper delivers additional insights, including a detailed comparative analysis of the participating algorithms and a discussion of the design choices and performance trade-offs for future method development.
@@
Sections 2 and  3 summarise the state of the art methods and provide an overview of trackerless freehand ultrasound reconstruction. Section 4 details the parameters of the Challenge, including the dataset, evaluation metrics and other Challenge setups. Section 5 describes the participation statistics and the methodologies submitted by participating teams, accompanied by the performance analysis of each method. Section 6 discusses the limitations of the Challenge and outlines potential directions for future work. Finally, Section 7 concludes the study by summarising the outcomes of the Challenge and highlighting its key contributions, benefits, and future directions.
2 Related Work
The field of 3D freehand ultrasound reconstruction has undergone a significant evolution over the past two decades. Early trackerless methods relied on statistical modeling to approximate probe motion without external tracking systems, such as speckle decorrelation analysis [1] and linear regression-based motion estimation [2]. While these approaches established the feasibility of trackerless scanning, they were often limited in robustness and generalisability, particularly in the presence of non-linear motion or varying anatomical structures [3, 4].
Recent advances in deep learning have led to a new wave of data-driven methods for trackerless 3D ultrasound reconstruction. Early CNN-based approaches [5] demonstrated that learned image representations could outperform traditional speckle-based techniques in estimating inter-frame motion. Subsequent work expanded on this by incorporating spatial and temporal modeling through Long Short-Term Memory (LSTM) networks [6, 7, 8, 9], and attention [10, 11], enabling more stable trajectory estimation over longer scan sequences. Transformer-based architectures [12, 13] have recently been introduced to better model long-range dependencies and spatial coherence.
Recent research in freehand 3D ultrasound reconstruction has pursued multiple directions to improve motion estimation accuracy, reconstruction speed, and generalisability across protocols and anatomies. One of the promising direction is the fusion of deep learning-based trajectory estimation with volumetric consistency optimisation. For instance, several methods combine learning-based inter-frame pose estimation with multi-view or global model refinement to improve alignment and robustness across large scan sequences [14, 15]. Another active area focuses on addressing domain shifts introduced by different ultrasound transducers. Domain adaptation strategies have been proposed to improve cross-device generalisation, particularly through the learning of transducer-invariant features [16]. A notable recent development involves the application of implicit neural representations to ultrasound imaging. These methods aim to compress the volumetric information and encode it as parameters of a model, offering potential gains in memory efficiency and spatial resolution compared with explicit representation [17, 18, 19]. [20] further extends this concept by adapting Gaussian Splatting techniques to volumetric ultrasound by replacing projection-based rendering with a model that aligns with ultrasound wave propagation. By leveraging anisotropic Gaussians, this method enables faster, memory efficient 3D ultrasound reconstruction.
Geometric and motion-aware modeling have also emerged as a powerful tool for improving reconstruction accuracy. Recent approaches integrate optical flow and spatiotemporal attention to better capture dense inter-frame motion and global spatiotemporal consistency [21, 22], while point cloud-based registration [23] techniques offer motion-driven alternatives to traditional pose estimation pipelines. In parallel, several works have investigated sensor signals and auxiliary tracking modalities to enhance pose estimation accuracy. IMUs have been integrated with learning-based systems to provide motion information or correct drift in trackerless trajectories [5, 24, 25, 26, 27]. Additionally, an evaluation of low-cost tracking alternatives has shown no statistically significant difference between high and low-end optical trackers, further supporting the feasibility of cost-effective freehand 3D ultrasound setups [28].
Despite these promising developments, several limitations persist. First, prior studies have typically relied on private or small-scale datasets, often with only 12-40 subjects, which limits both the statistical power of the findings and the generalisability of trained models. Second, the absence of standardised datasets and evaluation protocols has made it difficult to perform fair and reproducible comparisons between methods. As a result, reported performance varies widely across studies and may be influenced by dataset-specific biases or tuning.
In addition to the above-discussed practical limitations, technical difficulties are also found in other applied machine learning methods in medical imaging. The core lies in accurately estimating the 3D motion of the ultrasound probe from 2D image sequences alone. Unlike tracker-based systems, trackerless methods must compute probe motion from intensity patterns in the images. However, these patterns can be affected by factors that complicate motion estimation, such as tissue deformation due to varying probe pressure and changes in scanning speed or angle. This leads to accumulated drift over time, particularly in long or complex scanning trajectories, which directly impacts reconstruction accuracy and downstream clinical applications especially those requiring spatially accurate 3D volumes.
To date, few effort has been made on large-scale public benchmarking for trackless freehand ultrasound reconstruction. This Challenge was established to fill this gap by offering a comprehensive platform for evaluating trackerless freehand ultrasound reconstruction methods, providing a fair and rigorous environment for performance comparison. In summary, while significant progress has been made in trackerless ultrasound reconstruction, consistent benchmarking is lacking. This Challenge aims to establish a foundation for reproducible and scalable evaluation for trackerless freehand ultrasound reconstruction.
@@ 3 Preliminaries
The goal of freehand ultrasound reconstruction is to estimate the transformation between pairs of ultrasound frames within a scan without relying on any external tracking device, thereby enabling the reconstruction of 2D ultrasound images into 3D space.
@@ 3.1 Coordinate Systems and Spatial Transformations
Table 1 summarises the terminologies commonly used in freehand ultrasound reconstruction. In learning-based algorithm development, a tracking system is typically used to directly capture the transformation of each ultrasound frame, providing ground truth transformations for training supervision and for evaluating trackerless algorithms. The most commonly utilised tracking modalities are optical tracking systems and EM tracking systems. The optical tracking system consists of a camera that captures pose information and a rigid tracking tool attached to the ultrasound probe [29]. This tool typically includes at least three passive or active markers, which enable the determination of the probe’s six-degree-of-freedom (6-DoF) pose. After spatial calibration, as detailed in Section 3.2, the system can obtain the transformation matrix corresponding to each ultrasound frame. The tracking data are subsequently transferred and stored using an interface such as the open-source PLUS platform [30]. While ultrasound machine and tracking device typically have their own API for data access, tools such as PLUS simplify the process by offering a unified interface and consistent data and file formats, making practical integration more convenient, though not strictly necessary. The EM tracking system [31] comprises three main components: the transmitter, the system control unit, and the tracked receiver. When the probe is moved within the magnetic field produced by the transmitter, the receiver mounted on the probe detects induced electrical currents. The remainder of this section describes the three coordinate systems involved in freehand ultrasound reconstruction, using an optical tracking system as a representative example.
@@
There are three coordinate systems, as shown in Fig. 1(a): the image coordinate system, the tracker tool coordinate system, and the camera (or world) coordinate system. The image coordinate system defines the positions of pixels (in 2D) or voxels (in 3D) within an image. In the context of freehand ultrasound reconstruction, the image coordinate system specifically refers to the 2D coordinate system of individual ultrasound image frame. The transformation recorded by the optical camera is from tracker tool coordinate system to camera coordinate system, which in turn represents the location of the tracker in camera coordinate system. However, this tracker-reported transformation does not directly provide transformation to the coordinate system of the ultrasound image itself, from the other two coordinate systems. Consequently, a transformation, commonly referred to as the calibration matrix, is necessary to map the ultrasound image coordinates to the tracker tool coordinate system. This transformation is crucial for converting each pixel in the 2D ultrasound image to its corresponding voxel in the reconstructed 3D volume. The calibration matrix defines the transformation between the image coordinate system and the tracker tool coordinate system. It incorporates both a scaling factor that converts image coordinates from pixels to millimeters, as well as the spatial calibration that establishes the transformation between the image coordinate system (in millimeters) and the tracker tool coordinate system (in millimeters).
@@
Table 1: Terminologies in freehand ultrasound reconstruction.
@@
Tracker tool coordinate system A 3D coordinate system defined by three or four sphere markers which are attached to a rigid body with a unique geometry Origin of the marker attached to the object of interest (phantom, cadaver, patient, etc.) mm As defined by the tracking system / marker manufacturer
Camera (or world) coordinate system A 3D coordinate system defined by the tracking system manufacturer Origin of the tracking system (midpoint between the two camera lenses) mm X axis: increasing downward from the center between the two lenses; Y axis: increasing toward the camera’s right; Z axis: inward, toward the back of the device.
Terminology Definition
@@
$T_{scale}$ The transformation from the image coordinate system (in pixels) to the image coordinate system (in millimeters).
$T_{rotation}$ The transformation from the image coordinate system (in millimeters) to the tracker tool coordinate system (in millimeters).
@@
Let $T_{i}^{camera\leftarrow tool}$ denote the transformation matrix of frame $i$, recorded by the optical camera, representing the transformation from the tracker tool coordinate system of frame $i$ to the camera coordinate system. Then, the rigid transformation from image coordinate system (in millimeters) of frame $i$ to image coordinate system (in millimeters) of frame $j$, $T_{j\leftarrow i}$, is given by:
$T_{j\leftarrow i}=T_{rotation}^{-1}\cdot T_{j\leftarrow i}^{tool}\cdot T_{rotation}$ (1)
where $T_{j\leftarrow i}^{tool}$ denotes the transformation from the $i^{th}$ tracker tool to the $j^{th}$ tracker tool. $T_{rotation}=\left[\begin{array}[]{cc}\mathbf{R_{3\times 3}}&\mathbf{t_{3\times 1}}\\ \mathbf{0}&1\end{array}\right]$ represents the spatial calibration between the image coordinate system (in millimeters) and the tracker tool coordinate system, where $\mathbf{R}$ is a $3\times 3$ rotation matrix and $\mathbf{t}$ is a $3\times 1$ translation vector. The transformation $T_{rotation}$ is obtained through the calibration process, as described in Section. 3.2, while $T_{j\leftarrow i}^{tool}$ can be computed using Eq. (2).
@@
Reconstructing the 3D ultrasound volume and the trajectory of the ultrasound frames require determining the position of each frame. Let the first frame serve as the reference frame. If the transformations from each frame to the reference frame are known, the coordinates of all pixels within the scan can be computed using Eq. (3).
@@
where $p_{(u,v)}=(u,v,0,1)^{T}$ and $P_{(x,y,z)}=(x,y,z,1)^{T}$ represent the coordinates of pixel $(u,v)$ in the image coordinate system of the $i^{th}$ frame (in pixels) and the image coordinate system of the first frame (in millimeters), respectively. $T_{scale}=diag(s_{x},s_{y},1,1)$ is the scaling factor that converts from pixels to millimeters, represented by a diagonal matrix with elements $s_{x}$,$s_{y}$,$1$,$1$ along the main diagonal. $s_{x}$ and $s_{y}$ represent the scaling factor along the $x$ and $y$ axes respectively. The two $1$s represent no scaling on the $z$-axis and for the homogeneous coordinate.
When the distance between two ultrasound frames is too large, predicting the transformation becomes challenging. Therefore, a common approach is to estimate the transformation between two adjacent frames and then accumulate these transformations to obtain the desired result. The transformation from the $i^{th}$ frame to the first frame, $T_{1\leftarrow i}$, can be computed by recursively multiplying the previously estimated relative transformations, as shown in Eq. (4).
@@
Moreover, Eq. (4) illustrates that estimation errors can propagate and accumulate along the chain, ultimately leading to trajectory drift. To mitigate this issue, [7] leverages the long-term dependencies within the ultrasound sequence to estimate the transformations between frames at intermediate intervals.

### v2

@@ Abstract
Trackerless freehand ultrasound reconstruction aims to reconstruct 3D volumes from sequences of 2D ultrasound images without relying on external tracking systems. By eliminating the need for optical or electromagnetic trackers, this approach offers a low-cost, portable, and widely deployable alternative to more expensive volumetric ultrasound imaging systems, particularly valuable in resource-constrained clinical settings. However, predicting long-distance transformations and handling complex probe trajectories remain challenging. The TUS-REC2024 Challenge establishes the first benchmark for trackerless 3D freehand ultrasound reconstruction by providing a large publicly available dataset, along with a baseline model and a rigorous evaluation framework. By the submission deadline, the Challenge had attracted 43 registered teams, of which 6 teams submitted 21 valid dockerized solutions. The submitted methods span a wide range of approaches, including the state space model, the recurrent model, the registration-driven volume refinement, the attention mechanism, and the physics-informed model. This paper provides a comprehensive background introduction and literature review in the field, presents an overview of the challenge design and dataset, and offers a comparative analysis of submitted methods across multiple evaluation metrics. These analyses highlight both the progress and the current limitations of state-of-the-art approaches in this domain and provide insights for future research directions. All data and code are publicly available to facilitate ongoing development and reproducibility. As a live and evolving benchmark, it is designed to be continuously iterated and improved. The Challenge was held at MICCAI 2024 and is organised again at MICCAI 2025, reflecting its sustained commitment to advancing this field.
@@
TUS-REC2024 , Trackerless Freehand Ultrasound , 3D Reconstruction , Deep Learning , Spatial Transformation Estimation , MICCAI Challenge
@@
[queen_marry] organization=Centre for Bioengineering, Digital Environment Research Institute, School of Engineering and Materials Science, Queen Mary University of London, city=London, postcode=E1 4NS, country=U.K.
[…]
@@
[…]
[MUSIC_Lab_5] organization=School of Artificial Intelligence, Shenzhen University, city=Shenzhen, postcode=518060, country=China
@@ 1 Introduction
Ultrasound imaging remains a cost-effective, non-invasive modality with real-time capabilities, making it a valuable tool across a wide range of clinical applications [1]. However, it only provides incomplete 3D information because the locations of ultrasound frames are unknown. This poses challenges for applications requiring accurate volumetric information, such as biometric quantification, image registration, and 3D visualisation [2]. While expert clinicians can often infer 3D structure mentally or through standardised acquisition protocols (e.g., standard planes), the absence of inter-frame positional data limits reproducibility and the integration of ultrasound images into advanced image analysis workflows.
@@
Ongoing works seek to address this limitation by using 3D ultrasound probes to enable 3D reconstruction. 3D ultrasound probes are capable of acquiring volumetric data directly, using dedicated mechanical probes or 2D array transducers [3]. While these probes provide valuable 3D imaging capabilities and offer flexible scanning trajectories, their higher cost, limited user experience in clinical practice, and physical constraints (e.g., interference in 2D array transducers) may restrict their use in some clinical settings, such as low-resource environments, point-of-care scenarios, or mobile and emergency units where portability and affordability are critical [4].
@@
In comparison, freehand 2D ultrasound imaging has advantage of widespread availability and long-standing integration into clinical workflows. It has been used for decades across a broad range of applications, and clinicians are highly familiar with both the use and interpretation of it. Building on this established foundation, tracker-based freehand ultrasound reconstruction techniques have been introduced to enable the generation of 3D anatomical representations. These methods aim to enhance conventional 2D ultrasound by incorporating spatial information from external tracking systems, such as optical [5] or electromagnetic (EM) [6] trackers. This enables conventional 2D ultrasound probes to be used for 3D imaging, providing a more flexible and accessible solution in clinical and research applications where dedicated and bulky 3D ultrasound systems may be impractical. However, optical and EM tracking systems present other challenges in clinical environments. Optical tracking requires an unobstructed line of sight between the tracker and the camera [7]. Although approaches such as using multiple cameras have been proposed to mitigate this limitation, challenges related to system complexity and calibration remain [8, 9]. EM tracking is sensitive to nearby metal objects and electromagnetic interference, which can affect accuracy [10].
@@
Trackerless freehand ultrasound reconstruction refers to generating 3D volumetric representations from sequential 2D ultrasound frames in a handheld freehand scan, without using external tracking systems. This kind of method computes the relative spatial transformations among frames using images themselves. Common approaches include non-learning-based methods such as speckle decorrelation [11, 12, 13] and learning-based motion estimation such as convolutional neural network (CNN) [14, 15] and recurrent neural network (RNN) [16, 17]. Additionally, trackerless freehand ultrasound reconstruction may further enhance existing 3D ultrasound systems, rather than serving solely as alternatives.
@@
This application involves practical challenges such as handling both 2D and 3D imaging data, incorporating tracking information, and managing multiple spatial coordinate systems, all of which contribute to a significant barrier for newcomers to this field and may impede broader progress and adoption. In addition, trackless reconstruction encounters other challenges: 1) the difficulty of predicting accurate poses when the ultrasound sequence length is large; and 2) the high variability across different datasets, which complicates the validation and fair comparison of methods. While benchmarking may be essential to address this variability, progress has been limited by the scarcity of publicly available datasets, which are critical for both performance evaluation and the development of learning-based approaches.
@@
Furthermore, comparison of methods in the existing literature is often conducted on relatively small, private datasets, using a variety of evaluation metrics to assess performance [2]. This variability complicates the comparison of different methods’ strengths and weaknesses and may lead to biased conclusions, due to dataset characteristics, evaluation metric choices, and inherent differences in the methods’ underlying assumptions. For example, learning-based approaches may assume that training and testing data come from similar distributions, while classical methods may rely on consistently available speckle patterns.
@@
To bridge these gaps, we organised the TUS-REC2024 (Trackerless 3D Freehand Ultrasound Reconstruction) Challenge. This Challenge is designed to foster both algorithmic innovation and practical clinical applicability by promoting reproducibility, benchmarking, and methodological transparency. TUS-REC2024 Challenge provides an in vivo ultrasound dataset, consisting of scans from both the left and right forearms of 85 volunteers (2,040 scans, 1,025,448 frames in total), acquired using a time-synchronised optical tracking system. We aim to conduct a comprehensive comparison among methods, evaluating their strengths and weaknesses on a common dataset, using a consistent set of carefully-defined performance metrics. This approach will ensure a more objective and transparent assessment of the methods’ relative efficacy, and more importantly, to drive the development of new techniques for trackerless freehand ultrasound reconstruction.
@@
This Challenge has three key contributions. Firstly, it establishes a rigorous and standardised benchmarking framework for trackerless freehand ultrasound reconstruction, advancing the development of novel algorithms and promoting objective performance evaluation through withheld test data and unified assessment metrics. Secondly, it provides the necessary infrastructure to support this benchmarking effort, including a large publicly available dataset in the field, and detailed preliminary materials and accompanying code that describe the end-to-end pipeline for trackerless freehand ultrasound reconstruction. Thirdly, beyond the outcomes of the Challenge itself, this summary paper delivers additional insights, including a comprehensive literature review, a detailed comparative analysis of the participating algorithms and a discussion of the design choices and performance trade-offs for future method development.
@@
The rest of the article is organised as below. Sections 2 introduces the basic knowledge of trackerless freehand ultrasound reconstruction, aiming to provide researchers with clear technical background and consistent terminologies in this field. Sections 3 summarises the state-of-the-art methods in trackerless freehand ultrasound reconstruction, comparing traditional and deep learning techniques across key methodological aspects. Section 4 reports details of the Challenge, including dataset curation, evaluation metrics and so on. Section 5 describes the participation statistics and the methodologies submitted by participating teams, accompanied by performance analysis of each method. Section 6 discusses the limitations of the Challenge and outlines potential directions for future work. Finally, Section 7 concludes the study by summarising the outcomes of the Challenge and highlighting its key contributions, benefits, and future directions.
@@ 2 Preliminaries
The goal of trackerless freehand ultrasound reconstruction is to estimate the transformation between pairs of ultrasound frames within a scan without relying on any external tracking device, thereby enabling the reconstruction of 2D ultrasound images into 3D space. Table 1 summarises the terminologies commonly used in this field and the notations used across this paper.
@@ 2.1 Coordinate Systems and Spatial Transformations
For learning-based methods, a tracking system is typically used to directly capture the pose of each ultrasound frame, providing labels for training and ground truth for evaluation. The most commonly utilised tracking modalities are optical tracking systems and EM tracking systems. The optical tracking system consists of rigid tracking tools attached to the ultrasound probe and cameras that capture position of the tracking tool [18]. The tool typically includes at least three passive or active markers, which enable the determination of the probe’s six-degree-of-freedom (6-DoF) pose. After spatial calibration, as detailed in Section 2.2, the system can obtain the pose of each ultrasound frame itself. The tracking data are timestamped, and subsequently transferred and stored using an interface such as the open-source PLUS platform [19]. While ultrasound machine and tracking device typically have their own API for data management, softwares such as PLUS provide unified interface and consistent data formats, making practical integration more convenient, though not strictly necessary. The EM tracking system [20] comprises three main components: the transmitter, the system control unit, and the tracked receiver. When the probe is moved within the magnetic field produced by the transmitter, the receiver mounted on the probe detects induced electrical currents, for calculating the spatial location relative to a predefined reference.
The remainder of this section describes the three coordinate systems and their spatial transformations involved in freehand ultrasound reconstruction, using an optical tracking system as a representative example.
@@
As shown in Fig. 1(a), there are three coordinate systems: the image coordinate system, the tracker tool coordinate system, and the camera (or world) coordinate system, as defined in Table 1. The transformation recorded by the optical tracker is from tracker tool coordinate system to camera coordinate system, which represents the location of the tracking tool in camera coordinate system. However, this tracker-reported transformation does not directly provide transformation between the coordinate system of the ultrasound image itself and the other two coordinate systems. Consequently, a transformation, commonly referred to as the calibration matrix, is necessary to map the ultrasound image coordinate system to the tracker tool coordinate system. It usually incorporates both scaling matrix that converts image coordinate unit from pixels to millimeters, as well as rigid transformation between the image coordinate system (in millimeters) and the tracker tool coordinate system (in millimeters). This transformation is crucial for locating each pixel in the 2D ultrasound image in the 3D space for reconstruction purpose.
@@
Table 1: Terminologies in freehand ultrasound reconstruction and notations in this paper.
@@
Tracker tool coordinate system A 3D coordinate system defined by three or four sphere markers which are attached to a rigid body with a unique geometry Origin of the marker attached to the object of interest (phantom, cadaver, patient, etc.) $m$ As defined by the tracking system / marker manufacturer
Camera (or world) coordinate system A 3D coordinate system defined by the tracking system manufacturer Origin of the tracking system (midpoint between the two camera lenses) $m$ X axis: increasing downward from the center between the two lenses; Y axis: increasing toward the camera’s right; Z axis: inward, toward the back of the device.
Notation Definition
@@
$T_{scale}$ The scaling matrix to change the unit of the image coordinate system from pixels to millimeters.
$T_{rigid}$ The transformation from the image coordinate system (in millimeters) to the tracker tool coordinate system (in millimeters).
@@
Let $T_{i}^{camera\leftarrow tool}$ denote the transformation from the tracker tool coordinate system of frame $i$ to the camera coordinate system, as recorded by the optical tracking system during scanning. Then, the rigid transformation from image coordinate system (in millimeters) of frame $i$ to image coordinate system (in millimeters) of frame $j$, $T_{j\leftarrow i}$, is given by:
$T_{j\leftarrow i}=T_{rigid}^{-1}\cdot T_{j\leftarrow i}^{tool}\cdot T_{rigid}$ (1)
where $T_{j\leftarrow i}^{tool}$ denotes the transformation from the $i^{th}$ tracker tool to the $j^{th}$ tracker tool. $T_{rigid}=\left[\begin{array}[]{cc}\mathbf{R_{3\times 3}}&\mathbf{t_{3\times 1}}\\ \mathbf{0}&1\end{array}\right]$ represents the transformation from the image coordinate system (in millimeters) to the tracker tool coordinate system, where $\mathbf{R}$ is a $3\times 3$ rotation matrix and $\mathbf{t}$ is a $3\times 1$ translation vector. The transformation $T_{rigid}$ is obtained through the calibration process, as described in Section. 2.2, while $T_{j\leftarrow i}^{tool}$ can be computed using Eq. (2).
@@
Reconstructing the 3D ultrasound volume requires determining the position of each frame. Let the first frame serve as the reference frame. If the transformations from each frame to the reference frame are known, the coordinates of all pixels within the scan can be computed using Eq. (3).
@@
where $p_{(u,v)}=(u,v,0,1)^{T}$ and $P_{(x,y,z)}=(x,y,z,1)^{T}$ represent the homogeneous coordinates of the same point in the image coordinate system of the $i^{th}$ frame (in pixels) and the image coordinate system of the first frame (in millimeters), respectively. $T_{scale}=diag(s_{x},s_{y},1,1)$ is the diagonal scaling matrix that converts from pixels to millimeters, where $s_{x}$ and $s_{y}$ represent the scaling factors along the $x$ and $y$ axes respectively.
The transformation from the $i^{th}$ frame to the first frame, $T_{1\leftarrow i}$, can be computed by recursively multiplying the previously estimated relative transformations, as shown in Eq. (4).
@@
However, if any transformation in Eq. (4) has a large error, it may propagate and affect the subsequent trajectory.